OpenStudy (anonymous):

can someone help? The measures of two supplementary angles are in the ratio 2:7 what are the degree measures of the two angles?

OpenStudy (anonymous):

2/7 = x / (180 - x) 360 - 2x = 7x Can you finish it from here?

OpenStudy (anonymous):

Sorry I really dont understand much of it. From there do I simplify the equation?

OpenStudy (anonymous):

@tcarroll010

OpenStudy (anonymous):

mn + xn = (m + x)n

OpenStudy (anonymous):

what does m and n represent? haha..

OpenStudy (anonymous):

mn - xn = (m - x)n Use these distributive properties. Solve for "x" in my original equation. Angles are "x" and 180 - x

OpenStudy (anonymous):

"n" is anything you want it to be. It's a variable.

OpenStudy (anonymous):

Oh okay, let me try it one sec.

OpenStudy (anonymous):

I'm sorry I dont get it :( could you give me an outline or something?

OpenStudy (anonymous):

If you are having trouble with: 360 - 2x = 7x add 2x to both sides and get a new equation, one with a "numeric" on one side and on the other side you will have "x" with a coefficient in front of it. With this new equation, divide both sides by the coefficient.

OpenStudy (anonymous):

This start of adding 2x to both sides is done to get "x" on one side of the equation.

OpenStudy (anonymous):

Think of "x" as something tangible like apples. If you have 2 apples in one room and 7 in another, and you want to get all your apples together, you take the 2 apples from one of the rooms and put them with the rest in that other room. You will still have the same number of apples and that's the key concept.

OpenStudy (anonymous):

oh haha ok so it would be, 360-2x=7x 360=9x 360/9=40?

OpenStudy (anonymous):

x=40?

OpenStudy (anonymous):

There! You got it! and the other angle is 180 - x

OpenStudy (anonymous):

ok so the first angle is 40, and the second is 140?

OpenStudy (anonymous):

Notice that 40/140 is the same ratio as 2/7

OpenStudy (anonymous):

OH! thank you that helps alot! so when I write it all out I just use the equation then?

OpenStudy (anonymous):

And supplementary angles add to 180. That's why one is x and the other is: 180 - x x + (180 - x) = 180

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Ok thank you very much

OpenStudy (anonymous):

uw! Good luck to you in all of your studies and thx for the recognition! @baybay88